Write the standard form of the line that passes through the given points (6, 1) and (5, 4).
Answer by AlgebraLady88(44)

The standard form of a linear equation is
Ax + By = C
To find out what the equation looks like, we would first have to find the slope.
The formula for slope is y2 - y1 / x2 -x1
We will designate, from the points given x1 =6 y1= 1
x2 =5 y2= 4
So, slope is 4-1/5-6 = 3/-1 = -3
We would need to write the equation in slope-intercept form first: y=mx+b
To do that , we would need to find the y intercept , b . We will use the slope and one point . You can pick either (6,1) or (5,4)
We will use (6,1)
y= mx+ b
y= -3x + b
1= -3*6 + b
1= -18 +b
19= b
So, the slope intercept form would be y= -3x + 19
To change this to standard form :
y = -3x +19
3x+y = 19
To check , just plug in both (6,1) and (5,4) into the equation.
3x + y = 19
3(6) + 1 = 19
3x +y = 19
3 (5) + 4 = 19